OpenStudy (anonymous):

how to find a set of vectors reciprocal to the set : 3i - j + 2k , 2i + j -k and i - 2j + k

OpenStudy (anonymous):

let us say A = 3i-j+2k, B= 2i+j-k ,C=i-2j+k then reciprocal vector of A = 2*pi*( B cross C )/A . ( B cross C )

OpenStudy (anonymous):

similarly for B = 2 * pi * ( C cross A) / A.( B cross C) and C = 2 * pi * ( A cross B)/ A.(B cross C)

OpenStudy (anonymous):

thank you